OpenStudy (anonymous):

WILL GIVE MEDAL use binomial theorem to expand the binomial (d-4b)^3

OpenStudy (anonymous):

@hartnn help pls lol

OpenStudy (amistre64):

state the binomial thrm ... its assuming you have some prior knowledge for the question

OpenStudy (anonymous):

i dont understand it though, i understand the Pascal Tri, but not this

OpenStudy (amistre64):

the pascal triangle is just a visual representation of the binomial thrm ... a set of coefficients

OpenStudy (amistre64):

\[(a+b)^n=\sum_{k=0}^{n}\binom{n}{k}a^{n-k}~b^{k}\] this is a compact notation for a sum of a lot of terms

OpenStudy (anonymous):

okay im lost

OpenStudy (amistre64):

if we use the expanded version of it ... \[(a+b)^n=\binom{n}{0}a^{n}~b^{0}+\binom{n}{1}a^{n-1}~b^{1}+\binom{n}{2}a^{n-2}~b^{2}+...+\binom{n}{n}a^{n-n}~b^{n}\]

OpenStudy (amistre64):

what are you lost on?

OpenStudy (anonymous):

everything

OpenStudy (amistre64):

can you at least fill in the parts? they give you n,a, and b parts

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-64b^3+48b^2d-12bd^2+d^3

OpenStudy (anonymous):

would that be it?

OpenStudy (amistre64):

maybe .. do you have to do this by the binomial setup? or do they just expect a final result?

OpenStudy (anonymous):

i have to do the binomial setup

OpenStudy (anonymous):

ik the answer, but i want to know how to do it

OpenStudy (anonymous):

@Pleasehelpasap12345 will help you in a second, msgs wont work

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

let a = d let b = -4b let n = 3 \[(a+b)^n=\binom{n}{0}a^{n}~b^{0}+\binom{n}{1}a^{n-1}~b^{1}+\binom{n}{2}a^{n- 2}~b^{2}+...+\binom{n}{n}a^{n-n}~b^{n}\] \[(d+(-4b))^3=\binom{3}{0}d^{3}~(-4b)^{0}+\binom{3}{1}d^{3-1}~(-4b)^1\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\binom{3}{2}d^{3-2}~(-4b)^{2}+\binom{3}{3}d^{3-3}~(-4b)^{3}\] \[(d+(-4b))^3=\binom{3}{0}d^{3}-\binom{3}{1}4d^{2}b+\binom{3}{2}16db^{2}-\binom{3}{3}64b^{3}\] the parantheticals just relate to the pascal row values \[(d+(-4b))^3=(1)d^{3}-(3)4d^{2}b+(3)16db^{2}-(1)64b^{3}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still pretty confusing

OpenStudy (amistre64):

thats all it is ... filling in the parts good luckl